# Wave-particle correlations in multiphoton resonances of coherent light-matter interaction

### 07 Mar 2024 v2

@@ V Particles triggering the sampling of waves
Increasing the drive amplitude we enter the regime of steady-state bimodality where $g^{(2)}_{\rm ss}(0)<1$. All but one of the seven photon emissions visible in Fig. 5(b) are associated with jumps downwards in $\langle a^{\dagger}a(t)\rangle_{\rm REC}$. After the initial interruption of the beat, the semiclassical oscillation converges to the two steady-state attractors, while quantum jumps bring it close to the unstable state near the phase-space origin. Contrary to what we expect from steady-state antibunching, aligning the LO in the orthogonal direction ($\theta=\pi/4$) to the development of bimodality, generated a sample trajectory with a single closely-spaced pair, stopping the relaxation to the steady state with $\langle a^{\dagger}a\rangle_{\rm ss}\approx 0.54$. At the same time, the unconditional cross-correlation of the intensity and the field amplitude of the forwards-scattered light, $h_{\theta}(\tau)=\langle A_{\theta}(\tau)\rangle_{c}\equiv\langle a(\tau)\rangle_{c}e^{-i\theta}+{\rm c.c.}\,(\tau\geq 0)$, where $\langle a\rangle_{c}\equiv{\rm tr}[a\,e^{\mathcal{L}\tau}(a\rho_{\rm ss}a^{\dagger})]$, and its time-reversed version for $\tau\leq 0$ Carmichael et al. (2004), reveal an asymmetry of fluctuations Denisov et al. (2002) depending on the chosen value for $\theta$. An asymmetric $h_{\theta}(\tau)$ rules out the validity of Gaussian statistics and signals the breakdown of detailed balance Denisov et al. (2002) in the region of bimodality. A pronounced temporal asymmetry of intensity-field correlations before and after a trigger “click” has been identified for three-level atoms in Ref. Xu et al. (2015). For photon numbers $\sim 10^{-4}$ where we note the presence of strong bunching ($r=1$Shamailov et al. (2010) and a negative spectrum of squeezing ($r=0$), the correlation function attains very large negative values, violating its classical bounds as a result of the anomalous phase of the amplitude oscillation first noted for squeezed light in Carmichael et al. (2000).
@@
Figure 6: Wave-particle correlator unraveling the ME in a direction orthogonal to steady-state bi(tri)modality. (Top) Individual trajectory of the conditioned intracavity photon number (positive-valued plot in blue) and the conditioned field (plot with alternating sign in orange) forming part of the homodyne photocurrent [see Eq. (12)] for $r=1/2$. The operation point is the same as in Fig. 5(a), except for the phase of the local oscillator which is set to $\theta=\pi/4$. Inset A depicts the unlabeled spectrum of squeezing $\mathcal{S}_{\rightarrow}^{\theta}(\omega)$ against $(\omega-\omega_{0})/\kappa$ (upper plot) on top of the transmission spectrum $T(\omega)$ (lower plot) calculated from Eq. (43) [the four peaks labeled $a,b,c,d$ correspond to transitions $|\xi_{3}\rangle\to|\xi_{2}\rangle$, $|\xi_{1}\rangle\to|\xi_{0}\rangle$, $|\xi_{3}\rangle\to|\xi_{1}\rangle$ and $|\xi_{1}\rangle\to|\xi_{0}\rangle$, respectively], while inset B depicts a contour plot of the steady-state Wigner function. (Bottom) Schematic representation of the conditioned Wigner function (two projections) of the intracavity field at the time $\kappa t_{\rm max}\approx 9.95$ when the conditional photon number plotted in Fig. 5(d) reaches its maximum, $\langle a^{\dagger}a\rangle_{\rm REC,max}\approx 2.74$. The zero level is colored in green, midway between the main peak (yellow) and dip (blue).
@@
Furthermore, a few background comments are in order concerning the fluctuating conditioned field amplitude $\langle\psi_{c}(t)|A_{\theta}|\psi_{c}(t)\rangle$ plotted in all frames of Fig. 5, in conjunction with the photon number. This quantity is of particular importance since it forms part of the measured differential photocurrent together with the Gaussian white noise. There is a clear time asymmetry with respect to the origin set by each trigger photodetection at the APD. The BHD back action tends to recover the symmetry, albeit in a background of intense fluctuations. The back action comes about because of the quantum superposition of the local oscillator and the JC source field. These two fields should not be thought of as being separable since each time a local oscillator photon is detected the evolution of the cavity field is affected. This entails that the conditioned state $|\psi_{c}(t)\rangle$ at time $t$ is correlated with the shot noise at the detectors 1 and 2 from the recent past. Consequently, through the APD trigger clicks – the particle aspect in the duality – we postselect a subensemble of the shot noise which has been filtered through the dynamical response function of the source in the course of its coherent evolution.
@@
As a further instance of the wave-particle unraveling (with $r=1/2$), the top frame of Fig. 6 depicts an individual realization corresponding to a negative steady-state spectrum of squeezing (one obtained for $r=0$). Despite the presence of photon bunching [with $g_{\rm ss}^{(2)}(0)=2.17$ – see Fig. 4(b)] the three photon emissions are spaced further apart than the inverse of the average photon lifetime $(2\kappa)^{-1}$. Interestingly, the same number of emissions occurs compared to the unraveling with $\theta=3\pi/4$ where the quadrature amplitude is larger [see Fig. 5(a)]. The steady-state cavity photon number calculated from the solution of the ME (1) is $\langle a^{\dagger}a\rangle_{\rm ss}\approx 0.16$. In conjunction with the photon emission probabilities, the conditioned Wigner function depicted in the bottom frame of Fig. 6 – corresponding to the maximum conditional emission probability plotted in Fig. 5(a) – brings two aspects of the dynamics together: Quantum interference between dressed states produces a succession of peaks and dips with alternating signs as remnants of steady-state bistability, aligned in the direction imposed by the local oscillator (set at $\theta=0$) – once again, variations on the theme of conditionally generating single(few)-photon-added coherent states we met in Fig. 2(a). In the steady-state distribution computed from the time-averaged density operator (the time average is equivalent to the ensemble average since quantum trajectories are ergodic Carmichael (2008)) the interference is interrupted at random times, dephasing to the purely positive profiles of Fig. 5.
@@
where $A_{\theta}\equiv\frac{1}{2}(ae^{-i\theta}+a^{\dagger}e^{i\theta})$ is the quadrature amplitude selected by the local oscillator. From Eq. (15) we find that assuming Gaussian statistics for a weak source field, the wave-particle correlation of the fluctuations $h_{\theta}(\tau)-1$ [see Eq. (15)] and the spectrum of squeezing form a Fourier-transform pair, which means that small violations of nonclassicality in one part lead to large violations in the other (see Carmichael et al. (2000) and Sec. 2 of Carmichael et al. (2004)). On top of calculating unconditional quantum averages, we examine the regression of the fluctuations to steady state in the course of single trajectories generated in the wave-particle correlator unraveling. Note that the quantity $\langle:\Delta A_{\theta}(0)\Delta A_{\theta}(\tau):\rangle_{\rm ss}$ appearing in Eq. (15) is equal to $R_{\theta}(\tau)$ introduced in Sec. IV, since $\langle a\rangle_{\rm ss}=0$. As low strengths of the drive operationally qualify as those for which the detected forwards photon emission stream is highly bunched Shamailov et al. (2010).
@@ Appendix A Derivation of the waiting-time distribution
To assess the corpuscular nature of the forwards-scattered light, we employ the waiting-time distribution defined by conditional exclusive probability densities. This quantity is given by Carmichael et al. (1989); Car (1993)
@@
Finally, expressing the photon number operator in terms of the dressed states in the truncated Hilbert space as Shamailov et al. (2010); Mavrogordatos and Lledó (2021)
@@ Appendix B Derivation of the squeezing spectrum
with $\tilde{s}_{ij}=-i[(\omega-\omega_{0}+\tilde{E}_{j}/\hbar-\tilde{E}_{i}/\hbar)]/\kappa$, while the spectrum of squeezing in the forwards direction reads

### 13 Mar 2024 v3

@@ V Particles triggering the sampling of waves
Increasing the drive amplitude we enter the regime of steady-state bimodality where $g^{(2)}_{\rm ss}(0)<1$. All but one of the seven photon emissions visible in Fig. 5(b) are associated with jumps downwards in $\langle a^{\dagger}a(t)\rangle_{\rm REC}$. After the initial interruption of the beat, the semiclassical oscillation converges to the two steady-state attractors, while quantum jumps bring it close to the unstable state near the phase-space origin. Contrary to what we expect from steady-state antibunching, aligning the LO in the orthogonal direction ($\theta=\pi/4$) to the development of bimodality, generated a sample trajectory with a single closely-spaced pair, stopping the relaxation to the steady state with $\langle a^{\dagger}a\rangle_{\rm ss}\approx 0.54$. At the same time, the unconditional cross-correlation of the intensity and the field amplitude of the forward-scattered light, $h_{\theta}(\tau)=\langle A_{\theta}(\tau)\rangle_{c}\equiv\langle a(\tau)\rangle_{c}e^{-i\theta}+{\rm c.c.}\,(\tau\geq 0)$, where $\langle a\rangle_{c}\equiv{\rm tr}[a\,e^{\mathcal{L}\tau}(a\rho_{\rm ss}a^{\dagger})]$, and its time-reversed version for $\tau\leq 0$ Carmichael et al. (2004), reveal an asymmetry of fluctuations Denisov et al. (2002) depending on the chosen value for $\theta$. An asymmetric $h_{\theta}(\tau)$ rules out the validity of Gaussian statistics and signals the breakdown of detailed balance Denisov et al. (2002) in the region of bimodality. A pronounced temporal asymmetry of intensity-field correlations before and after a trigger “click” has been identified for three-level atoms in Ref. Xu et al. (2015). For photon numbers $\sim 10^{-4}$ where we note the presence of strong bunching ($r=1$Shamailov et al. (2010) and a negative spectrum of squeezing ($r=0$), the correlation function attains very large negative values, violating its classical bounds as a result of the anomalous phase of the amplitude oscillation first noted for squeezed light in Carmichael et al. (2000).
@@
Figure 6: Wave-particle correlator unraveling the ME in a direction orthogonal to steady-state bi(tri)modality. (Top) Individual trajectory of the conditioned intracavity photon number (positive-valued plot in blue) and the conditioned field (plot with alternating sign in orange) forming part of the homodyne photocurrent [see Eq. (12)] for $r=1/2$. The operation point is the same as in Fig. 5(a), except for the phase of the LO which is set to $\theta=\pi/4$. Inset A depicts the unlabeled spectrum of squeezing $\mathcal{S}_{\rightarrow}^{\theta}(\omega)$ against $(\omega-\omega_{0})/\kappa$ (upper plot) on top of the transmission spectrum $T(\omega)$ (lower plot) calculated from Eq. (43) [the four peaks labeled $a,b,c,d$ correspond to transitions $|\xi_{3}\rangle\to|\xi_{2}\rangle$, $|\xi_{1}\rangle\to|\xi_{0}\rangle$, $|\xi_{3}\rangle\to|\xi_{1}\rangle$ and $|\xi_{1}\rangle\to|\xi_{0}\rangle$, respectively], while inset B depicts a contour plot of the steady-state Wigner function. (Bottom) Schematic representation of the conditioned Wigner function (two projections) of the intracavity field at the time $\kappa t_{\rm max}\approx 9.95$ when the conditional photon number plotted in Fig. 5(d) reaches its maximum, $\langle a^{\dagger}a\rangle_{\rm REC,max}\approx 2.74$. The zero level is colored in green, midway between the main peak (yellow) and dip (blue).
@@
Furthermore, a few background comments are in order concerning the fluctuating conditioned field amplitude $\langle\psi_{c}(t)|A_{\theta}|\psi_{c}(t)\rangle$ plotted in all frames of Fig. 5, in conjunction with the photon number. This quantity is of particular importance since it forms part of the measured differential photocurrent together with the Gaussian white noise. There is a clear time asymmetry with respect to the origin set by each trigger photodetection at the APD. The BHD back action tends to recover the symmetry, albeit in a background of intense fluctuations. The back action comes about because of the quantum superposition of the LO and the JC source field. These two fields should not be thought of as being separable since each time a LO photon is detected the evolution of the cavity field is affected. This entails that the conditioned state $|\psi_{c}(t)\rangle$ at time $t$ is correlated with the shot noise at the detectors 1 and 2 from the recent past. Consequently, through the APD trigger clicks – the particle aspect in the duality – we postselect a subensemble of the shot noise which has been filtered through the dynamical response function of the source in the course of its coherent evolution.
@@
As a further instance of the wave-particle unraveling (with $r=1/2$), the top frame of Fig. 6 depicts an individual realization corresponding to a negative steady-state spectrum of squeezing (one obtained for $r=0$). Despite the presence of photon bunching [with $g_{\rm ss}^{(2)}(0)=2.17$ – see Fig. 4(b)] the three photon emissions are spaced further apart than the inverse of the average photon lifetime $(2\kappa)^{-1}$. Interestingly, the same number of emissions occurs compared to the unraveling with $\theta=3\pi/4$ where the quadrature amplitude is larger [see Fig. 5(a)]. The steady-state cavity photon number calculated from the solution of the ME (1) is $\langle a^{\dagger}a\rangle_{\rm ss}\approx 0.16$. In conjunction with the photon emission probabilities, the conditioned Wigner function depicted in the bottom frame of Fig. 6 – corresponding to the maximum conditional emission probability plotted in Fig. 5(a) – brings two aspects of the dynamics together: Quantum interference between dressed states produces a succession of peaks and dips with alternating signs as remnants of steady-state bistability, aligned in the direction imposed by the LO (set at $\theta=0$) – once again, variations on the theme of conditionally generating single(few)-photon-added coherent states we met in Fig. 2(a). In the steady-state distribution computed from the time-averaged density operator (the time average is equivalent to the ensemble average since quantum trajectories are ergodic Carmichael (2008)) the interference is interrupted at random times, dephasing to the purely positive profiles of Fig. 5.
@@
where $\Delta A_{\theta}\equiv\frac{1}{2}(\Delta ae^{-i\theta}+\Delta a^{\dagger}e^{i\theta})$ is the quadrature amplitude fluctuation operator probed by the setting of the LO phase. From Eq. (15) we find that assuming Gaussian statistics for a weak source field, the wave-particle correlation of the fluctuations $h_{\theta}(\tau)-1$ [see Eq. (15)] and the spectrum of squeezing form a Fourier-transform pair, which means that small violations of nonclassicality in one part lead to large violations in the other (see Carmichael et al. (2000) and Sec. 2 of Carmichael et al. (2004)). On top of calculating unconditional quantum averages, we examine the regression of the fluctuations to steady state in the course of single trajectories generated in the wave-particle correlator unraveling. Note that the quantity $\langle:\Delta A_{\theta}(0)\Delta A_{\theta}(\tau):\rangle_{\rm ss}$ appearing in Eq. (15) is equal to $R_{\theta}(\tau)$ introduced in Sec. IV, since $\langle a\rangle_{\rm ss}=0$. As low strengths of the drive operationally qualify as those for which the detected forward photon emission stream is highly bunched Shamailov et al. (2010).
@@ Appendix A Derivation of the waiting-time distribution
To assess the corpuscular nature of the forward-scattered light, we employ the waiting-time distribution defined by conditional exclusive probability densities. This quantity is given by Carmichael et al. (1989); Car (1993)
@@
Finally, we express the photon-number operator in terms of the dressed states in the truncated Hilbert space as Shamailov et al. (2010); Mavrogordatos and Lledó (2021)
@@ Appendix B Derivation of the squeezing spectrum
with $\tilde{s}_{ij}=-i[(\omega-\omega_{0}+\tilde{E}_{j}/\hbar-\tilde{E}_{i}/\hbar)]/\kappa$, while the spectrum of squeezing in the forward direction reads